OpenStudy (anonymous):

prakash ,sameer, ramesh and tarun have total 240 rs with them. Prakash has a half the total amount of what the others have. sameer has one-third of the total amount of what the others have. Ramesh has one-fourth of the total amount of what the others have.find the amount with tarun in rs?

terenzreignz (terenzreignz):

I hate systems of linear equations... BUT... we do what we must ^_^

terenzreignz (terenzreignz):

We should probably represent what they have as variables. Let's get started :D Prakash = p sameer = s ramesh = r tarun = t

terenzreignz (terenzreignz):

Now, @nachi1991 based on what the problem says, what is p + s + r + t = ?

OpenStudy (anonymous):

amount with each of them equals 240..right?

terenzreignz (terenzreignz):

That's right. p + r + s + t = 240 four unknowns... we need exactly four equations... so we need three more.

terenzreignz (terenzreignz):

Can you work out another equation?

terenzreignz (terenzreignz):

Use this statement: "Prakash has a half the total amount of what the others have. "

OpenStudy (anonymous):

p=240/2...???

terenzreignz (terenzreignz):

No... Let's break it down... after I demonstrate this, the other two will be easy :) "Prakash has a half the total amount of what the others have" "Prakash has..."\[\Large \color{red}p=\] "...half..." \[\Large p = \frac12[]\] "...the total amount of what the OTHERS have" \[\Large p = \frac12[\color{red}{r+s+t}]\] Does that make sense to you? ^_^

terenzreignz (terenzreignz):

r + s + t is the total amount that the others have ^_^

OpenStudy (anonymous):

ya k...so sameer will have s=1/3(p+r+t)..is it ok?

terenzreignz (terenzreignz):

That's right :) And what about Ramesh?

OpenStudy (anonymous):

R=1/4(p+s+t)..right?..

terenzreignz (terenzreignz):

Yes :) You now have your four equations... \[\Large p+r+s+t = 240\]\[\Large p = \frac12[r+s+t]\]\[\Large s= \frac13[p+r+t]\]\[\Large r = \frac14[p+s+t]\] That's four equations... ^_^

terenzreignz (terenzreignz):

I can see a creative way to solve this... if you're interested :) Because I think... solving this in traditional substitution or elimination is too long and cumbersome... right @uri? XD

terenzreignz (terenzreignz):

Okay, you ready? :)

OpenStudy (anonymous):

now what do i do with these equations?

terenzreignz (terenzreignz):

Let's look at these two equations first... \[\Large p+r+s+t=240\]and\[\Large p = \frac12[r+s+t]\] okay?

OpenStudy (anonymous):

ya it will be great if u have some creative way to solve it..

terenzreignz (terenzreignz):

Okay, look at this... \[\Large p = \frac12[r+s+t]\] What happens if you multiply both sides by 2?

terenzreignz (terenzreignz):

Hey @nachi1991 work with me here ^_^

terenzreignz (terenzreignz):

Okay, if you multiply both sides by 2, you get \[\Large 2p = r+s+t\] right? since 2 and 1/2 cancel out...

terenzreignz (terenzreignz):

Focus :P @nachi1991 Does this now make sense to you? 2p = r + s + t

OpenStudy (anonymous):

ya i got it..terenzreignz..

OpenStudy (anonymous):

hey i m sorry bt this uri is disturbing me a lot.

terenzreignz (terenzreignz):

You can call me TJ. And so, we can replace the r+s+t here: \[\Large p + \color{red}{r+s+t}= 240\] with 2p, since they are equal: \[\Large p + \color{red}{2p}= 240\] And now, you can solve for p easily ^_^

OpenStudy (anonymous):

k TJ yes we can replace so it will become 3p=240 i.e p=80 right?

terenzreignz (terenzreignz):

That's right :D Okay keep that in mind, p = 80. but we won't use it. Instead, we'll look at the third equation: \[\Large s = \frac13[p+r+t] \]multiply both sides by 3, we get?

OpenStudy (anonymous):

3s=p+r+t yes?

terenzreignz (terenzreignz):

Indeed :) So, we have a p + r + t here in the first equation: \[\Large \color{red}{p+r}+s +\color{red}t=240\] which we can replace with 3s since they are equal... \[\Large \color{red}{3s}+s = 240\] solve for s? :P

OpenStudy (anonymous):

ya k so it will be s=60.

terenzreignz (terenzreignz):

Correct :) Okay, p = 80, s = 60 Now we'll use the fourth equation: \[\Large r = \frac14[p+s+t]\] You know the drill... multiply both sides by 4, you get?

terenzreignz (terenzreignz):

Sorry, I'm in a bit of a rush, so... 4r = p + s + t Now, back to the first equation again... \[\Large \color{red}p +r +\color{red}{s+t}=240 \] p+s+t becomes 4r because of equality... \[\Large \color{red}{4r}+r = 240\] so now solve for r.

OpenStudy (anonymous):

ya so r=48....and then we have to put values of p r and right?...oh..u r awesome..thank u so much...

terenzreignz (terenzreignz):

Yes. We know that p = 80, s = 60 and r = 48. So using the first equation, you can now easily solve for t. Do it... I have to go now :D \[\Large p +s +r +t = 240\]\[\Large 80 + 60 + 48 + t = 240\] I'm sure you can wrap it up now, and solve for the last variable (t) Signing off now --------------------------------- Terence out ^_^

OpenStudy (anonymous):

good bye TJ..thnx